OpenStudy (anonymous):

The amount of interest earned on savings is directionally proportional to the amount of money saved. If $87.50 interest is earned on $2,500 of savings, how much interest will be earned on $3,500 of savings over the same time period? A. $105.00 B. $122.50 C. $140.00 D. $157.50

OpenStudy (texaschic101):

87.50 / 2500 = x/3500 --- (87.50 to 2500 = x to 3500) cross multiply 2500x = (3500)(87.50) 2500x = 306250 x = 306250/2500 x = 122.50 (B)